1.The aggregate production function is Y = F(K, L) where Y = real GDP, K = capital and L = labor. For Y = F(K, L) = K1/3Lb where b = 0 there are

A.decreasing returns to scale B.increasing returns to scale C.constant returns to scale D.none of the other answers is correct. E. we cannot determine the returns to scale

2.If y = f(x) = ln U(x) then dy/dx = f’(x) =

a. none of the other answers is correct. b.e^U(x)

C.U’(x)/U(x) D. e^U'(x) E. U’(x)
Answer #1

oggregate Production function 3 Y = F (K16)= K b L where b=0 => y = f(K 16)= k wl- L -M (2)=1 =) Y= F (KIC)= K (1) Y = F (k,(2) y-flal- in u(x) du dx U(X) 0 l_ UK) dy 0(x) dx (X) Ane: (C) 0(x) O(x)for quetsion (2)

First we take derivative of ln(×)= 1/x then the derivative of x